# for Scalable MoE Pruning

### v1

@@ 7 Conclusion
We probe $(\lambda_{1},\lambda_{2})\in\{(0,1),(1,0),(1,1)\}$, except for the Snowflake Arctic, which is the biggest MoE we deal with, where we only consider $(\lambda_{1},\lambda_{2})=(1,0)$, which means no GPU calls is needed for expert pruning. To get coactivation values $a_{i,j}$, we utilize 1000 samples from the C4 dataset, each of which has 2048 sequence length. We evaluate on lm-evaluation-harness Gao et al. (2021) We use 4bit quantization Dettmers et al. (2023) for experiments with Mixtral-8x22B and Arctic, due to their model size. We use 20%, 12.5%, and 10% for Arctic, Mixtral-8x7B, and Mixtral-8x22B respectively, as the expert pruning ratio for STUN. These are the maximum values among 10, 12.5, 20, 25, and 35%, with minimum performance loss. We use $\kappa=3$ for selective reconstruction. For Wanda and OWL, we use 128 C4 samples following the original papers Yin et al. (2024); Sun et al. (2024), while we use 4096 for sequence length. For OWL, we use the default setting, $M=5,\lambda=0.08$.
@@
Other Clustering Algorithms
@@
where $t_{DSatur}$ is some threshold to control the pruning ratio of MoE.
Ablation Studies Results
Detailed per-task performance of ablation studies are described in Tables 4, 5.

### v2

@@ Appendix B Implementation Details
We probe $(\lambda_{1},\lambda_{2})\in\{(0,1),(1,0),(1,1)\}$, except for the Snowflake Arctic, which is the biggest MoE we deal with, where we only consider $(\lambda_{1},\lambda_{2})=(1,0)$, which means no GPU calls is needed for expert pruning. To get coactivation values $a_{i,j}$, we utilize 1000 samples from the C4 dataset, each of which has 2048 sequence length. We evaluate on lm-evaluation-harness Gao et al. (2021) We use 4bit quantization Dettmers et al. (2023) for experiments with Mixtral-8x22B and Arctic, due to their model size. We use 20%, 12.5%, and 10% for Arctic, Mixtral-8x7B, and Mixtral-8x22B respectively, as the expert sparsity for STUN. These are the maximum values among 10, 12.5, 20, 25, and 35%, with minimum performance loss. We use $\kappa=3$ for selective reconstruction. For Wanda and OWL, we use 128 C4 samples following the original papers Yin et al. (2024); Sun et al. (2024), while we use 4096 for sequence length. For OWL, we use the default setting, $M=5,\lambda=0.08$.
@@
Appendix C Other Clustering Algorithms
@@
where $t_{DSatur}$ is some threshold to control the sparsity of MoE.
Cluster Reconstruct LM-eval Avg
Ours Ours 59.58
DSatur Ours 58.59
Ours Always 57.60
Ours Never 59.22
Appendix D Ablation Studies
To validate our design of expert pruning in sections A.1 and A.2, we evaluate alternative approaches to expert-prune Mixtral-8x7B at 50% sparsity. Table 8 confirms that our design choices are valid. Our agglomerative clustering algorithm outperforms the DSatur algorithm, an alternative clustering algorithm we discuss in the Appendix. Additionally, selective reconstruction proves superior to always or never reconstructing, as shown in the last two rows. Detailed per-task performance of ablation studies are described in Tables 9, 10. Detailed hyperparameter ablation studies are described in Tables 11, 12. Results with applying expert-pruning only are provided in Table 14.
@@ Appendix D Ablation Studies
$\lambda_{1}$ $\lambda_{2}$ GSM8K
Mixtral-8x7B 1 0 63.38
0 1 58.53
1 1 60.42
Mixtral-8x22B 1 0 81.50
0 1 81.58
1 1 80.52
Table 11: Comparison of different $\lambda_{1}$, $\lambda_{2}$ configurations. For the Arctic, we only consider the cheapest, ($\lambda_{1}$, $\lambda_{2}$) = (1,0).
sparsity GSM8K
Arctic 10% 70.74
20% 69.90
35% 60.05
Mixtral-8x7B 12.50% 63.38
25% 53.22
Mixtral-22B 10% 81.58
12.50% 79.91
25% 74.07
Table 12: Comparison of different sparsity of our proposed expert-level pruning. ($\lambda_{1}$, $\lambda_{2}$) is probed among (1,0),(0,1),(1,1), except for 10% and 20% of Arctic.
25% Ours $O(1)$ 59.30 85.44 88.13 64.42 64.52 35.40 71.84 75.37 68.05
Lu et al. (2024a) $O(\frac{k^{n}}{\sqrt{n}})$ 58.19 84.89 87.34 65.24 62.47 35.60 70.04 75.85 67.45
25% Ours $O(1)$ 52.73 81.82 83.09 60.84 63.34 31.60 68.59 72.69 64.34
Lu et al. (2024a) $O(\frac{k^{n}}{\sqrt{n}})$ 51.62 81.94 83.64 61.60 58.72 33.00 67.87 75.37 64.22
sparsity GSM8K Avg (→) ARC-c ARC-e HellaSwag MMLU
Arctic 20% 69.90 67.95 57.00 83.71 65.70 65.39
Mixtral-8x7B 12.50% 63.38 70.65 61.18 86.66 66.80 67.97
Mixtral-8x22B 10% 81.58 71.78 61.18 85.82 66.62 73.52
Table 14: Performance when applying only our expert pruning.
Appendix E Detailed Results for RQ2
Table 13 describes the per-task performance of RQ2.
Appendix F Why GSM8K is Harder
The nature of the GSM8K task, which is a generation task, accounts for this discrepancy. A random generation baseline on GSM8K would achieve 0% accuracy, making it far more challenging to maintain performance. In contrast, ARC, HellaSwag, and MMLU are multiple-choice tasks where random baselines can achieve reasonable accuracy by comparing the perplexity of different completion options. As a result, maintaining performance on GSM8K is considerably harder, a challenge often overlooked in previous works. This explains the more pronounced difference in performance between STUN and the baselines on GSM8K.